# RL-U2Net: A Dual-Branch UNet with Reinforcement Learning-Assisted Multimodal Feature Fusion for Accurate 3D Whole-Heart Segmentation

### v1

@@ Abstract
CodePublic˙After˙Acceptance.
@@ Introduction
Cardiovascular disease (CVD) represents a leading cause of mortality worldwide. Accurate three-dimensional whole-heart segmentation is essential for quantitative lesion assessment and clinical decision-making. While computed tomography (CT) and magnetic resonance imaging (MRI) serve as primary diagnostic tools, single-modality approaches suffer from inherent limitations in contrast, spatial resolution, and imaging artifacts that compromise comprehensive cardiac characterization. Effective multi-modal feature fusion methods are therefore critical for enhancing segmentation accuracy and robustness in clinical applications (Valsangiacomo Buechel and Mertens 2012; Puyol-Antón et al. 2022).
@@
Deep learning has significantly advanced CVD diagnosis through medical imaging analysis. U-shaped architectures have become the dominant paradigm for medical segmentation due to their multi-scale feature aggregation and skip connections (Ronneberger, Fischer, and Brox 2015; Jin et al. 2020). However, traditional 3D CNNs suffer from limited receptive fields, hindering long-range dependency modeling (Çiçek et al. 2016). While Transformers’ self-attention mechanisms capture global correlations and achieve notable progress in visual segmentation (Carion et al. 2020; Strudel et al. 2021; Cao et al. 2022), they exhibit weaker fine-grained local feature representation and higher computational costs. To address these limitations, hybrid CNN-Transformer frameworks have emerged (Chen et al. 2021; Wang et al. 2022; Chen et al. 2021). This integration of CNN’s local discriminative capabilities with Transformer’s global dependency modeling has proven effective for improving cardiac segmentation accuracy.
Cardiac imaging presents unique challenges due to the heart’s non-rigid dynamic structure, where single-modality approaches often yield insufficient information. CT provides superior spatial resolution and vascular visualization, while MRI excels in soft tissue contrast and multi-parameter characterization (Freed et al. 2016). In recent years, multi-modal cardiac image segmentation has garnered significant attention. Researchers typically employ registration, fusion, or domain-adaptive methods to address data distribution discrepancies between different modalities (Li et al. 2023a), as shown in Figure 1. However, existing multi-modal segmentation methods face three critical limitations. First, substantial spatial misalignment occurs between modalities due to cardiac motion and respiratory variations, making traditional image-level registration inadequate for precise correspondence. Second, most fusion strategies rely on static concatenation or simple weighting, lacking deep understanding of inter-modal semantic relationships and limiting feature expressiveness. Third, many approaches decouple feature alignment from segmentation objectives, preventing end-to-end optimization and reducing overall performance.
@@
To overcome 3D CNN limitations in long-range dependency modeling, UNETR integrates Transformer encoders with U-shaped decoders (Hatamizadeh et al. 2022), while Swin-UNETR employs hierarchical shifted window attention with self-supervised pre-training (Tang et al. 2022; Hatamizadeh et al. 2021). Recent advances include axial global attention (GASA-UNet) (Sun et al. 2024), Mamba-based state space models (EM-Net) (Chang et al. 2024), and multi-scale convolution-attention fusion (Pan et al. 2025), reflecting efforts to balance computational efficiency with global context modeling. However, these methods primarily target single-modal scenarios and lack explicit handling of spatial inconsistencies and semantic correspondences in multimodal data, making feature alignment and fusion critical performance bottlenecks.
@@
Multimodal cardiac segmentation has attracted considerable research interest (Zhuang and Li 2020). Existing approaches fall into three categories: registration-based methods that spatially align modalities before segmentation (Luo and Zhuang 2022; Zhuang 2018; Luo and Zhuang 2020); fusion-based methods that exploit complementary CT-MRI characteristics through input-level (Yu et al. 2020; Zhang, Noga, and Punithakumar 2020), feature-level (Zhao, Boutry, and Puybareau 2020; Li et al. 2022), or decision-level fusion (Rokach 2010) with attention mechanisms; and domain adaptation methods using adversarial learning or style transfer for cross-domain generalization (Pei et al. 2021; Koehler et al. 2021; Wang and Zheng 2022). However, these methods typically treat registration and fusion as preprocessing steps (Li et al. 2023b), making unified end-to-end optimization of features alignment, fusion, and segmentation a persistent challenge.
@@
Reinforcement learning (RL) learns optimal policies through agent-environment interactions to maximize long-term rewards (Kaelbling, Littman, and Moore 1996). Deep RL employs neural networks for policy and value function modeling, enabling high-dimensional applications (Arulkumaran et al. 2017). In vision tasks, RL’s iterative ”perception-correction-feedback” process proves particularly effective for complex organ segmentation with ambiguous boundaries. Recent pixel-level RL methods have improved multi-organ boundary accuracy (Liu et al. 2025), while RL agents excel in image navigation, keypoint localization, and contour refinement (Alansary et al. 2018; Ghesu et al. 2016; Liao et al. 2020). Proximal Policy Optimization (PPO) achieves optimal balance among sample efficiency, stability, and implementation complexity through clipped surrogate objectives that constrain policy updates (Schulman et al. 2017). With the development of multimodal large models, PPO has been widely used for cross-modal policy optimization (Wan et al. 2025; Huang et al. 2025; He et al. 2016). Inspired by this, this paper innovatively introduces the PPO algorithm into medical image segmentation, utilizing reinforcement learning to assist in multimodal feature alignment.
@@
The core design of the PoseAlign module is based on the theory of cube rotational symmetry, constructing a complete action space comprising 24 rotational transformations (Worrall and Brostow 2018). These 24 rotations encompass all possible orientations of a cube in 3D space. The specific generation process is achieved by combining the six face orientations (±$\pm x,\pm y,\pm z$ axis directions) with four rotation angles (0°, 90°, 180°, 270°) for each face. The mathematical representation of the rotation matrix is:
=$\mathcal{R}=\{R_{1},R_{2},\ldots,R_{24}\}\subset SO(3)$ (6)
@@
The training process based on the PPO algorithm is the core driving mechanism of the RL-XAlign module, which models cross-modal feature alignment as a Markov decision process and implements collaborative learning of policy optimization and value assessment through the Actor-Critic architecture (Yao et al. 2024). This training strategy uses experience replay and policy pruning mechanisms to ensure the stability and convergence of the training process. The state space of the reinforcement learning environment is designed as a three-dimensional vector representation:
@@
$\mathcal{L}_{Policy}=-\mathbb{E}\left[A_{t}\cdot\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})},\right.\right.$
@@
Where $\mathcal{L}_{slow}$ and $\mathcal{L}_{fast}$ represent the loss values of the slow convergence mode and fast convergence mode, respectively, $\gamma$ is the temperature parameter controlling the sensitivity of weight adjustment, and $\delta$ is the baseline offset providing stability assurance. The selection of the hyperbolic tangent function ensures that the weight values vary within the (1,1)(-1,1)( - 1 , 1 ) interval, avoiding training instability caused by extreme weights. Based on the calculated weight factors, the balanced loss is reconstructed using an adaptive weighting strategy:
@@ Results
The dataset used in this study was obtained from the MM-WHS Challenge 2017 (Zhuang et al. 2019), which includes 60 sets of cardiac CT and MRI image data. Following the data partitioning strategy adopted in previous studies (Cui et al. 2023, 2025), the 40 sets of data were used as the training set, while the 20 sets of labeled data were randomly divided into 15 sets for testing and 5 sets for validation. The rest of the information of dataset is included in the supplementary materials.
@@
3D U-Net (Çiçek et al. 2016) 0.894 0.909 0.917 0.869 0.891 0.933 0.883 0.899 22.988
ConResNet (Lee et al. 2022) 0.918 0.929 0.928 0.883 0.914 0.949 0.852 0.910 26.652
nnformer (Zhou et al. 2023) 0.866 0.916 0.923 0.899 0.917 0.935 0.873 0.904 12.174
D-Former (Wu et al. 2023) 0.860 0.892 0.918 0.903 0.920 0.937 0.886 0.902 14.760
@@
(Hatamizadeh et al. 2021)
UNETR++ (Shaker et al. 2024) 0.883 0.881 0.924 0.899 0.893 0.934 0.860 0.896 14.850
@@
(Salgado-Garcia et al. 2024)
HRMedSeg (Xu et al. 2025) 0.910 0.924 0.937 0.913 0.920 0.951 0.892 0.921 12.255
P2TC (Cui et al. 2025) 0.907 0.930 0.936 0.894 0.918 0.953 0.889 0.918 21.417
@@
To comprehensively evaluate the segmentation performance of RL-U2Net, we systematically compared it with eleven state-of-the-art segmentation models on the MM-WHS 2017 dataset, as shown in the Table 1. To ensure fairness in the comparison, all comparison methods were re-experimented locally using the same dataset and evaluation metrics, with Dice coefficient and Hausdorff distance as the primary evaluation metrics. Training strategy and implementation details is included in the supplementary materials.
In CT image segmentation tasks, RL-U2Net demonstrated outstanding overall performance, with an average Dice coefficient of 93.1%, which is 1 percentage point higher than the second-best method, HRMedSeg (Xu et al. 2025), with 92.1%. In terms of boundary accuracy, RL-U2Net had an average Hausdorff distance of 11.471 mm, which was significantly better than all other methods. Results of MRI image segmentation tasks is included in the supplementary materials.
@@
To thoroughly validate the effectiveness of each key component of RL-U2Net, we designed a series of ablation experiments to assess the contribution of each core module to the overall segmentation performance by removing them one by one (see Table 2 for details).
@@ Appendix A Supplementary Contents
Due to visual limitation, in the main text, we only give a brief description of the backbone network. In order to show the structure of the network and related modules more clearly, we provide a more detailed description of the overall architecture of RL-U2Net in the supplementary material. The overall architecture of the proposed RL-U2Net is shown in Figure 6. The main network of RL-U2Net consists of a shared encoder based on Swin Transformer, an RL-XAlign cross-modal alignment module, a Res-Fusion fusion module, and a dual-branch ResU-Net decoder. For the input CT and MRI images $I_{\rm{CT}},I_{\rm{MRI}}\in\mathbb{R}^{H\times W\times D}$, the data first undergoes standardization and enhancement through a data preprocessing module, followed by patch segmentation to convert the 3D images into overlapping patch sequences as the network’s input. The encoder adopts a hierarchical design consisting of four Swin Transformer stages (Figure 7a)), with an RL-XAlign module embedded after each stage to achieve cross-modal feature alignment. The Swin Transformer module is responsible for extracting local-global multi-scale feature representations, while the RL-XAlign module achieves cross-modal feature alignment to address spatial inconsistency issues. At the end of each encoder stage, a Patch Merging operation is performed for downsampling, halving the spatial resolution of the feature map while doubling the channel dimension, thereby progressively constructing a hierarchical feature representation. Inspired by the U-Net architecture, we designed two independent and symmetric ResBlock decoder branches (He et al. 2016) (Figure 7b)), corresponding to the segmentation tasks for CT and MRI modalities, respectively. Each decoder consists of five upsampling stages, which restore the feature map resolution through deconvolution layers. Each upsampling doubles the spatial size while halving the number of channels. Jump connections are established between the encoder and decoder, and the Res-Fusion module (see Figure 7c)) effectively fuses the high-resolution shallow features of the encoder with the high-semantic deep features of the decoder to compensate for the loss of detail information during the downsampling process. Finally, the ensemble learning decision module generates precise whole-heart segmentation results through sliding window inference and weighted voting mechanisms. The following sections will detail the design principles and implementation mechanisms of each core module.
@@
[…]
@@
28:   $\mathcal{L}_{\text{critic}}\leftarrow\texttt{MSE}(V_{\text{now}},\hat{V})$
@@
3D U-Net (Çiçek et al. 2016) 0.686 0.844 0.873 0.824 0.808 0.783 0.755 0.796 39.996
ConResNet (Lee et al. 2022) 0.848 0.887 0.922 0.865 0.847 0.811 0.775 0.851 42.411
nnformer (Zhou et al. 2023) 0.682 0.814 0.848 0.846 0.824 0.787 0.763 0.795 31.557
D-Former (Wu et al. 2023) 0.732 0.849 0.887 0.891 0.849 0.811 0.784 0.829 31.097
@@
(Hatamizadeh et al. 2021)
UNETR++ (Shaker et al. 2024) 0.711 0.829 0.883 0.883 0.848 0.807 0.813 0.825 30.240
@@
(Salgado-Garcia et al. 2024)
HRMedSeg (Xu et al. 2025) 0.843 0.885 0.933 0.894 0.877 0.815 0.827 0.868 27.616
P2TC (Cui et al. 2025) 0.837 0.890 0.928 0.896 0.892 0.835 0.817 0.871 45.493
@@
A more detailed description of the dataset and data preprocessing methods are added in this section. The dataset used in this study was obtained from the MM-WHS Challenge 2017 (Zhuang et al. 2019) [49], which includes 60 sets of cardiac CT and MRI image data. Among these, 20 sets were manually annotated and used as training data, while the remaining 40 sets of unannotated data were used as test data. Since only 20 sets of publicly annotated data were available, the dataset needed to be re-divided to avoid overfitting. Following the data partitioning strategy adopted in previous studies (Cui et al. 2023, 2025), the 40 sets of data were used as the training set, while the 20 sets of labeled data were randomly divided into 15 sets for testing and 5 sets for validation. To ensure the consistency of the training data and the stable convergence of the model, all images underwent a standardized preprocessing workflow. This includes: resampling images to a uniform voxel spacing of (1.0,1.0,1.5)(1.0,1.0,1.5)( 1.0 , 1.0 , 1.5 ) mm to balance resampling quality and label integrity; Performing intensity normalization, where CT image intensity ranges are standardized to [175,250][-175,250][ - 175 , 250 ] and mapped to [0,1][0,1][ 0 , 1 ], and MRI image intensity ranges are standardized to [150,1488][150,1488][ 150 , 1488 ] and mapped to [0,1][0,1][ 0 , 1 ].
@@
Limitations and Future Work
Despite achieving state-of-the-art performance, our study still has some limitations. We have only evaluated it on the MM-WHS 2017 dataset, and external validation on more whole-heart segmentation benchmarks is lacking. Our approach fundamentally relies on paired multimodal data from the same subjects, where CT and MRI acquisitions capture the same anatomical structures with complementary imaging features, allowing for meaningful cross-modal feature alignment and information fusion. However, publicly available cardiac datasets with paired and annotated CT-MRI from the same subjects are still severely limited, thus limiting systematic testing of method generalisability. In future studies, we will further expand the size of the training and validation data to enhance the robustness of the model. And further extend this framework to be able to be applied to a wider range of medical image segmentation tasks.

### v2

@@ Abstract
Codehttps://github.com/TantalumKevin/RL-U2NET
@@ Introduction
Cardiovascular disease (CVD) represents a leading cause of mortality worldwide. Accurate three-dimensional whole-heart segmentation is essential for quantitative lesion assessment and clinical decision-making. While computed tomography (CT) and magnetic resonance imaging (MRI) serve as primary diagnostic tools, single-modality approaches suffer from inherent limitations in contrast, spatial resolution, and imaging artifacts that compromise comprehensive cardiac characterization. Effective multi-modal feature fusion methods are therefore critical for enhancing segmentation accuracy and robustness in clinical applications (valsangiacomobuechel2012imaging; puyol-anton2022multimodal).
@@
Deep learning has significantly advanced CVD diagnosis through medical imaging analysis. U-shaped architectures have become the dominant paradigm for medical segmentation due to their multi-scale feature aggregation and skip connections (ronneberger2015unet; jin2020raunet). However, traditional 3D CNNs suffer from limited receptive fields, hindering long-range dependency modeling (cicek20163d). While Transformers’ self-attention mechanisms capture global correlations and achieve notable progress in visual segmentation (carion2020endtoend; strudel2021segmenter; cao2022swinunet), they exhibit weaker fine-grained local feature representation and higher computational costs. To address these limitations, hybrid CNN-Transformer frameworks have emerged (chen2021transunet; wang2022mixed; chen2021transunet). This integration of CNN’s local discriminative capabilities with Transformer’s global dependency modeling has proven effective for improving cardiac segmentation accuracy.
Cardiac imaging is challenged by the heart’s non-rigid dynamics, often necessitating multi-modal approaches as single modalities provide incomplete information (freed2016mr). Consequently, multi-modal cardiac segmentation has drawn significant interest, typically employing registration, fusion, or domain adaptation to mitigate inter-modal discrepancies (li2023multimodality) (Figure 1). However, existing methods suffer from three critical limitations. First, substantial spatial misalignment, induced by cardiac and respiratory motion, renders traditional image-level registration inadequate for precise correspondence. Second, most fusion strategies rely on static concatenation or simple weighting, lacking deep understanding of inter-modal semantic relationships and limiting feature expressiveness. Third, many approaches decouple feature alignment from the segmentation objective, precluding end-to-end optimization and compromising overall performance.
@@
To overcome 3D CNN limitations in long-range dependency modeling, UNETR integrates Transformer encoders with U-shaped decoders (hatamizadeh2022unetr), while Swin-UNETR employs hierarchical shifted window attention with self-supervised pre-training (tang2022selfsupervised; hatamizadeh2021swin). Recent advances include axial global attention (GASA-UNet) (sun2024gasaunet), Mamba-based state space models (EM-Net) (chang2024emnet), and multi-scale convolution-attention fusion (pan2025multiscale), reflecting efforts to balance computational efficiency with global context modeling. However, these methods primarily target single-modal scenarios and lack explicit handling of spatial inconsistencies and semantic correspondences in multimodal data, making feature alignment and fusion critical performance bottlenecks.
@@
Multimodal cardiac segmentation has attracted considerable research interest (zhuang2020myocardial). Existing approaches fall into three categories: registration-based methods that spatially align modalities before segmentation (luo2022mathcal; zhuang2018multivariate; luo2020mvmmregnet); fusion-based methods that exploit complementary CT-MRI characteristics through input-level (yu2020dual; zhang2020fully), feature-level (zhao2020stacked; li2022myocardial), or decision-level fusion (rokach2010ensemblebased) with attention mechanisms; and domain adaptation methods using adversarial learning or style transfer for cross-domain generalization (pei2021disentangle; koehler2021unsupervised; wang2022cycmis). However, these methods typically treat registration and fusion as preprocessing steps (li2023myops), making unified end-to-end optimization of features alignment, fusion, and segmentation a persistent challenge.
@@
Reinforcement learning (RL) learns optimal policies through agent-environment interactions to maximize long-term rewards (kaelbling1996reinforcement). Deep RL employs neural networks for policy and value function modeling, enabling high-dimensional applications (arulkumaran2017brief). In vision tasks, RL’s iterative ”perception-correction-feedback” process proves particularly effective for complex organ segmentation with ambiguous boundaries. Recent pixel-level RL methods have improved multi-organ boundary accuracy (liu2025pixel), while RL agents excel in image navigation, keypoint localization, and contour refinement (alansary2018automatic; ghesu2016artificial; liao2020iterativelyrefined). Proximal Policy Optimization (PPO) achieves optimal balance among sample efficiency, stability, and implementation complexity through clipped surrogate objectives that constrain policy updates (schulman2017proximal). With the development of multimodal large models, PPO has been widely used for cross-modal policy optimization (wan2025srpo; huang2025highresolution; he2016deep). Inspired by this, this paper innovatively introduces the PPO algorithm into medical image segmentation, utilizing reinforcement learning to assist in multimodal feature alignment.
@@
The core design of the PoseAlign module is based on the theory of cube rotational symmetry, constructing a complete action space comprising 24 rotational transformations (worrall2018cubenet). These 24 rotations encompass all possible orientations of a cube in 3D space. The specific generation process is achieved by combining the six face orientations ($\pm x,\pm y,\pm z$ axis directions) with four rotation angles (0°, 90°, 180°, 270°) for each face. The mathematical representation of the rotation matrix is:
$\mathcal{R}=\{R_{1},R_{2},\ldots,R_{24}\}\subset SO(3)$ (6)
@@
The training process based on the PPO algorithm is the core driving mechanism of the RL-XAlign module, which models cross-modal feature alignment as a Markov decision process and implements collaborative learning of policy optimization and value assessment through the Actor-Critic architecture (yao2024cmragent). This training strategy uses experience replay and policy pruning mechanisms to ensure the stability and convergence of the training process. The state space of the reinforcement learning environment is designed as a three-dimensional vector representation:
@@
[…]
@@
Where $\mathcal{L}_{slow}$ and $\mathcal{L}_{fast}$ represent the loss values of the slow convergence mode and fast convergence mode, respectively, $\gamma$ is the temperature parameter controlling the sensitivity of weight adjustment, and $\delta$ is the baseline offset providing stability assurance. The selection of the hyperbolic tangent function ensures that the weight values vary within the (1,1)(-1,1) interval, avoiding training instability caused by extreme weights. Based on the calculated weight factors, the balanced loss is reconstructed using an adaptive weighting strategy:
@@ Results
The dataset used in this study was obtained from the MM-WHS Challenge 2017 (zhuang2019evaluation), which includes 60 sets of cardiac CT and MRI image data. Following the data partitioning strategy adopted in previous studies (cui2023improved; cui2025p2tc), the 40 sets of data were used as the training set, while the 20 sets of labeled data were randomly divided into 15 sets for testing and 5 sets for validation. The rest of the information of dataset is included in the supplementary materials.
@@
3D U-Net (cicek20163d) 0.894 0.909 0.917 0.869 0.891 0.933 0.883 0.899 22.988
ConResNet (lee20223d) 0.918 0.929 0.928 0.883 0.914 0.949 0.852 0.910 26.652
nnformer (zhou2023nnformer) 0.866 0.916 0.923 0.899 0.917 0.935 0.873 0.904 12.174
D-Former (wu2023dformer) 0.860 0.892 0.918 0.903 0.920 0.937 0.886 0.902 14.760
@@
UNETR++ (shaker2024unetr) 0.883 0.881 0.924 0.899 0.893 0.934 0.860 0.896 14.850
@@
(salgado-garcia2024efficient)
HRMedSeg (xu2025hrmedseg) 0.910 0.924 0.937 0.913 0.920 0.951 0.892 0.921 12.255
P2TC (cui2025p2tc) 0.907 0.930 0.936 0.894 0.918 0.953 0.889 0.918 21.417
@@
To comprehensively evaluate the segmentation performance of RL-U2Net, we systematically compared it with nine state-of-the-art segmentation models on the MM-WHS 2017 dataset, as shown in the Table 1. To ensure fairness in the comparison, all comparison methods were re-experimented locally using the same dataset and evaluation metrics, with Dice coefficient and Hausdorff distance as the primary evaluation metrics. Training strategy and implementation details is included in the supplementary materials.
In CT image segmentation tasks, RL-U2Net demonstrated outstanding overall performance, with an average Dice coefficient of 93.1%, which is 1 percentage point higher than the second-best method, HRMedSeg (xu2025hrmedseg), with 92.1%. In terms of boundary accuracy, RL-U2Net had an average Hausdorff distance of 11.471 mm, which was significantly better than all other methods. Results of MRI image segmentation tasks is included in the supplementary materials.
@@
To thoroughly validate the effectiveness of each key component of RL-U2Net, we designed a series of ablation experiments to assess the contribution of each core module to the overall segmentation performance by removing them one by one (see Table 2 for details). For more detailed ablation experiment contents, please refer to the appendix.
@@ Appendix A Supplementary Contents
Due to visual limitation, in the main text, we only give a brief description of the backbone network. In order to show the structure of the network and related modules more clearly, we provide a more detailed description of the overall architecture of RL-U2Net in the supplementary material. The overall architecture of the proposed RL-U2Net is shown in Figure 6. The main network of RL-U2Net consists of a shared encoder based on Swin Transformer, an RL-XAlign cross-modal alignment module, a Res-Fusion fusion module, and a dual-branch ResU-Net decoder. For the input CT and MRI images $I_{\rm{CT}},I_{\rm{MRI}}\in\mathbb{R}^{H\times W\times D}$, the data first undergoes standardization and enhancement through a data preprocessing module, followed by patch segmentation to convert the 3D images into overlapping patch sequences as the network’s input. The encoder adopts a hierarchical design consisting of four Swin Transformer stages (Figure 7a)), with an RL-XAlign module embedded after each stage to achieve cross-modal feature alignment. The Swin Transformer module is responsible for extracting local-global multi-scale feature representations, while the RL-XAlign module achieves cross-modal feature alignment to address spatial inconsistency issues. At the end of each encoder stage, a Patch Merging operation is performed for downsampling, halving the spatial resolution of the feature map while doubling the channel dimension, thereby progressively constructing a hierarchical feature representation. Inspired by the U-Net architecture, we designed two independent and symmetric ResBlock decoder branches (he2016deep) (Figure 7b)), corresponding to the segmentation tasks for CT and MRI modalities, respectively. Each decoder consists of five upsampling stages, which restore the feature map resolution through deconvolution layers. Each upsampling doubles the spatial size while halving the number of channels. Jump connections are established between the encoder and decoder, and the Res-Fusion module (see Figure 7c)) effectively fuses the high-resolution shallow features of the encoder with the high-semantic deep features of the decoder to compensate for the loss of detail information during the downsampling process. Finally, the ensemble learning decision module generates precise whole-heart segmentation results through sliding window inference and weighted voting mechanisms. The following sections will detail the design principles and implementation mechanisms of each core module.
@@
7:   $(d_{\text{now}},s_{\text{now}})\leftarrow\texttt{model.forward}(\texttt{ct},\texttt{mr})$
@@
[…]
@@
3D U-Net (cicek20163d) 0.686 0.844 0.873 0.824 0.808 0.783 0.755 0.796 39.996
ConResNet (lee20223d) 0.848 0.887 0.922 0.865 0.847 0.811 0.775 0.851 42.411
nnformer (zhou2023nnformer) 0.682 0.814 0.848 0.846 0.824 0.787 0.763 0.795 31.557
D-Former (wu2023dformer) 0.732 0.849 0.887 0.891 0.849 0.811 0.784 0.829 31.097
@@
UNETR++ (shaker2024unetr) 0.711 0.829 0.883 0.883 0.848 0.807 0.813 0.825 30.240
@@
(salgado-garcia2024efficient)
HRMedSeg (xu2025hrmedseg) 0.843 0.885 0.933 0.894 0.877 0.815 0.827 0.868 27.616
P2TC (cui2025p2tc) 0.837 0.890 0.928 0.896 0.892 0.835 0.817 0.871 45.493
@@
A more detailed description of the dataset and data preprocessing methods are added in this section. In this study, we mainly used the MM-WHS dataset to evaluate the performance of the model we proposed, and then conducted supplementary experiments using the MyoPS 2020 data to verify the generalization performance of the model.
The MM-WHS 2017 dataset used in this study was obtained from the MM-WHS Challenge 2017 (zhuang2019evaluation) [49], which includes 60 sets of cardiac CT and MRI image data. Among these, 20 sets were manually annotated and used as training data, while the remaining 40 sets of unannotated data were used as test data. Since only 20 sets of publicly annotated data were available, the dataset needed to be re-divided to avoid overfitting. Following the data partitioning strategy adopted in previous studies (cui2023improved; cui2025p2tc), the 40 sets of data were used as the training set, while the 20 sets of labeled data were randomly divided into 15 sets for testing and 5 sets for validation. To ensure the consistency of the training data and the stable convergence of the model, all images underwent a standardized preprocessing workflow. This includes: resampling images to a uniform voxel spacing of $(1.0,1.0,1.5)$ mm to balance resampling quality and label integrity; Performing intensity normalization, where CT image intensity ranges are standardized to [175,250][-175,250] and mapped to $[0,1]$, and MRI image intensity ranges are standardized to $[150,1488]$ and mapped to $[0,1]$.
The MyoPS dataset had 45 cases of multi-sequence CMR (25 cases for training and 20 cases for testing), each of which refers to a patient with three sequence CMR, i.e., LGE, T2 and bSSFP CMR. The data have been pre-processed using the MvMM method (zhuang2016multivariate; zhuang2018multivariate), to align the three-sequence CMR into a common space and to resample them into the same spatial resolution. The provided gold-standard labels include LV blood pool, RV blood pool, LV normal myocardium, LV myocardial edema, and LV myocardial scars. To focus on the segmentation of cardiac structure, we combined myocardium, myocardial edema and myocardial scar into a single myocardial category (yan2021novel).
@@
Results on MyoPS Dataset
We conducted supplementary experiments on the public MyoPS2020 dataset. This dataset includes three cardiac magnetic resonance (CMR) modalities: bSSFP, LGE, and T2-SPAIR. We implemented three cross-modal configurations—bSSFP+LGE, T2+LGE, and T2+bSSFP—to comprehensively assess the robustness of cross-sequence fusion. The results are shown in Table 4.
When bSSFP served as the primary modality, using LGE as auxiliary input yielded an average Dice of 85.73% and an HD95 of 6.92. This slightly outperformed the configuration using the T2 modality as auxiliary. When LGE was the primary modality, using bSSFP as auxiliary input achieved the best overall performance in this study, with an average Dice of 87.90% and an HD95 of 7.95. Finally, when T2-SPAIR was used as the main source and bSSFP was used as the auxiliary source to obtain Avg Dice=86.59% and HD95=8.18; The results of LGE as an auxiliary source were Avg Dice= 86.01% and HD95=9.53.
These supplementary results demonstrate that our proposed RL-U2Net framework achieves consistently robust and effective segmentation performance across diverse cross-modal combinations on the MyoPS2020 dataset. The bSSFP modality, in particular, demonstrated significant informational value, contributing to strong performance both as a primary and an auxiliary source.
Auxiliary Dice $\uparrow$ HD95 $\downarrow$
modality Myo LV RV Avg
bSSFP cine CMR results
LGE 83.06 85.49 88.63 85.73 6.92
T2 82.52 85.23 87.78 85.18 7.59
LGE CMR results
bSSFP 87.22 88.17 88.30 87.90 7.95
T2 85.84 87.95 87.37 87.05 11.49
T2-SPAIR CMR results
bSSFP 87.48 87.14 85.14 86.59 8.18
LGE 86.80 86.51 84.74 86.01 9.53
Table 4: Supplementary experiments on the MyoPS dataset.